# Evaluating Disentangled Representations for Controllable Music Generation

### v1

@@ 2.1 Models and Disentanglement Strategies
  • $S$-VQ-VAE [5], designed for style transfer, disentangles content and style through a discrete codebook for content and a style encoder regularized with data augmentations.
@@ 3.1 Model Training and Configurations
The three models $S$-VQ-VAE, TS-DSAE, and AFTER are trained using their official repositories [^1] For AFTER, we rely on the real-time model implementation. with default configurations, assumed to be optimal for each disentanglement strategy despite differing embedding sizes. In addition, we train two variants of AFTER to isolate specific disentanglement strategies: one without pitch and tempo augmentations applied to the timbre encoder input (AFTER-no-aug), and one without the adversarial loss (AFTER-no-adv). Table 1 summarizes the different model configurations under study.
@@ 4 Results
Informativeness. As shown in Table 2, $S$-VQ-VAE performs best on instrument classification for timbre, likely due to its larger embedding size. This aligns with [19], which links informativeness to embedding capacity, and the same trend extends to most structure tasks. Remarkably, it also achieves the best multi-pitch estimation performance despite its lower temporal resolution; however, overall performance on this note-level task remains low. In contrast, TS-DSAE is most effective for tempo regression, suggesting that its architecture and training strategy compensate for its smaller embedding size. AFTER remains less consistent, with notably weaker tempo prediction and only minor variant-specific effects.
Equivariance. Table 3 summarizes the equivariance results, revealing a negative correlation with informativeness and embedding size: $S$-VQ-VAE, while highly informative, is less equivariant. This trade-off is consistent with observations in the original framework [19]. In contrast, TS-DSAE, despite its compact structure embedding, achieves the strongest R-equivariance across all transformations. AFTER attains intermediate performance, with variants showing reduced timbre equivariance and mixed outcomes for structure.
Invariance. Table 4 presents the invariance results. Timbre embeddings are robust to tempo perturbations but sensitive to pitch shifts. $S$-VQ-VAE shows the strongest pitch invariance, followed by AFTER. Its variant without adversarial loss is most vulnerable, while removing pitch and tempo augmentations surprisingly results in higher invariance to both factors. Structure embeddings are broadly invariant to instrument changes, though AFTER again weakens when adversarial loss is removed. Consistent with the original framework [19], invariance appears driven less by capacity and more by training objectives and disentanglement strategies.
Disentanglement. Table 5 summarizes the disentanglement results. $S$-VQ-VAE shows little timbre leakage into the structure embedding, but its timbre embedding carries substantial structural information, likely due to its higher capacity and despite data augmentations. AFTER, in contrast, reveals the opposite trend: its structure embedding contains notable timbre information, while the timbre embedding generally carries little pitch information. Additional structural leakage is observed in variants without disentanglement strategies. Across models, timbre embeddings systematically encode tempo, as concatenation consistently improves tempo regression performance. Overall, TS-DSAE appears the most balanced, with concatenation having only marginal effects.

### v2

@@ 2.1 Models and Disentanglement Strategies
  • SS-VQ-VAE [6], designed for style transfer, disentangles content and style through a discrete codebook for content and a style encoder regularized with data augmentations.
@@ 3.1 Model Training and Configurations
The three models SS-VQ-VAE, TS-DSAE, and AFTER are trained using their official repositories [^1] For AFTER, we rely on the real-time model implementation. with default configurations, assumed to be optimal for each disentanglement strategy despite differing embedding sizes. In addition, we train two variants of AFTER to isolate specific disentanglement strategies: one without pitch and tempo augmentations applied to the timbre encoder input (AFTER-no-aug), and one without the adversarial loss (AFTER-no-adv). Table 1 summarizes the different model configurations under study.
@@ 4 Results
Informativeness. As shown in Table 2, SS-VQ-VAE performs best on instrument classification for timbre, likely due to its larger embedding size. This aligns with [20], which links informativeness to embedding capacity, and the same trend extends to most structure tasks. Remarkably, it also achieves the best multi-pitch estimation performance despite its lower temporal resolution; however, overall performance on this note-level task remains low. In contrast, TS-DSAE is most effective for tempo regression, suggesting that its architecture and training strategy compensate for its smaller embedding size. AFTER remains less consistent, with notably weaker tempo prediction and only minor variant-specific effects.
Equivariance. Table 3 summarizes the equivariance results, revealing a negative correlation with informativeness and embedding size: SS-VQ-VAE, while highly informative, is less equivariant. This trade-off is consistent with observations in the original framework [20]. In contrast, TS-DSAE, despite its compact structure embedding, achieves the strongest R-equivariance across all transformations. AFTER attains intermediate performance, with variants showing reduced timbre equivariance and mixed outcomes for structure.
Invariance. Table 4 presents the invariance results. Timbre embeddings are robust to tempo perturbations but sensitive to pitch shifts. SS-VQ-VAE shows the strongest pitch invariance, followed by AFTER. Its variant without adversarial loss is most vulnerable, while removing pitch and tempo augmentations surprisingly results in higher invariance to both factors. Structure embeddings are broadly invariant to instrument changes, though AFTER again weakens when adversarial loss is removed. Consistent with the original framework [20], invariance appears driven less by capacity and more by training objectives and disentanglement strategies.
Disentanglement. Table 5 summarizes the disentanglement results. SS-VQ-VAE shows little timbre leakage into the structure embedding, but its timbre embedding carries substantial structural information, likely due to its higher capacity and despite data augmentations. AFTER, in contrast, reveals the opposite trend: its structure embedding contains notable timbre information, while the timbre embedding generally carries little pitch information. Additional structural leakage is observed in variants without disentanglement strategies. Across models, timbre embeddings systematically encode tempo, as concatenation consistently improves tempo regression performance. Overall, TS-DSAE appears the most balanced, with concatenation having only marginal effects.